OpenStudy (anonymous):

Solve. –6 ≥ 10 – 8x (Points : 1) x ≥ 2 x ≥ –2 x ≤ 2 x ≤ –2

OpenStudy (anonymous):

@jim_thompson5910 oh and also thank you for not giving me the answer and making me figure it out because otherwise i would still be stuck do you think you could help me on this one also

OpenStudy (anonymous):

im giving another medal

jimthompson5910 (jim_thompson5910):

10 – 8x is the same as -8x + 10

jimthompson5910 (jim_thompson5910):

so –6 ≥ 10 – 8x is the same as –6 ≥ -8x + 10

OpenStudy (anonymous):

it would be -2 and u flip the sign right @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what you first do is subtract 10 from both sides –6 ≥ -8x + 10 –6 - 10 ≥ -8x + 10 - 10 -16 ≥ -8x what's next?

OpenStudy (anonymous):

do you divide 8 on both sides @jim_thompson5910

jimthompson5910 (jim_thompson5910):

-8, yes

jimthompson5910 (jim_thompson5910):

the sign will flip

OpenStudy (anonymous):

so it would be positive 2

OpenStudy (anonymous):

and you flip the sign

jimthompson5910 (jim_thompson5910):

so we will have \[\large 2 \le x\] and that turns into \[\large x \ge 2\]

OpenStudy (anonymous):

–5t ≥ 65 t ≤ –13 so is this the correct answer for this one

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (anonymous):

ok just checking

OpenStudy (anonymous):

so the sign never flipped because for the 2 the sign is the same as it was before so does it or does it not flip

jimthompson5910 (jim_thompson5910):

well it flips twice once when you divide both sides by a negative number then it does a second time when you flip the whole expression

jimthompson5910 (jim_thompson5910):

when I moved that x over I mean

OpenStudy (anonymous):

ohhhh okay so it did

OpenStudy (anonymous):

5 + d > 5 – d (Points : 1) d > 10 d > 0 d < –5 d > –10

OpenStudy (anonymous):

im confused on this one very @jim_thompson5910

OpenStudy (anonymous):

what do i do with the 5 where do i move it or change

jimthompson5910 (jim_thompson5910):

ok this will be the last one and then I have to go

jimthompson5910 (jim_thompson5910):

you subtract 5 from both sides

jimthompson5910 (jim_thompson5910):

and that will turn 5 + d > 5 - d into d > -d

OpenStudy (anonymous):

ok thnk u for helping

jimthompson5910 (jim_thompson5910):

then you add 'd' to both sides to go from d > -d to 2d > 0 and finally you divide both sides by 2 to get d > 0

OpenStudy (anonymous):

so would it be -5 or -10

OpenStudy (anonymous):

ohhhh so u dont have either number its just a 0

jimthompson5910 (jim_thompson5910):

yes exactly

jimthompson5910 (jim_thompson5910):

when you subtract 5 from both sides, they go away

OpenStudy (anonymous):

may i ask you one more that i had for last because i really dont get this one at all

jimthompson5910 (jim_thompson5910):

ok one more

OpenStudy (anonymous):

The sum of two consecutive integers is at least 46. What is the least possible pair of integers? (Points : 1) 21 and 23 23 and 24 22 and 23 24 and 25

jimthompson5910 (jim_thompson5910):

hint: "The sum of two consecutive integers is at least 46." translates to \[\large x + (x+1) \ge 46\] solve for x

OpenStudy (anonymous):

how do u know how to solve all of these lol do you think you could be my algebra tutor

jimthompson5910 (jim_thompson5910):

I could help from time to time, yes, or you can ask for help on this website and there are a lot of helpers here too. I know from lots of practice. Basically if x is any number, then x+1 is the next consecutive number so that's how I got x + (x+1) and this must be "at least 46" which means \[\large x + (x+1) \ge 46\]

OpenStudy (anonymous):

like do u have gmail kik or anything tht we can stay in contact on

jimthompson5910 (jim_thompson5910):

you can message me through here and that works fine I think

OpenStudy (anonymous):

and does tht mean the numbers tht add up to it or are the closest

jimthompson5910 (jim_thompson5910):

what do you mean?

OpenStudy (anonymous):

the question The sum of two consecutive integers is at least 46. What is the least possible pair of integers? (Points : 1) 21 and 23 23 and 24 22 and 23 24 and 25

OpenStudy (anonymous):

are u suppose to get the closest numbers or what can add up t this

jimthompson5910 (jim_thompson5910):

yeah you have to solve \[\large x + (x+1) \ge 46\] for x. then use this to find x+1

OpenStudy (anonymous):

so how do i do tht

OpenStudy (anonymous):

i mean how do i find x

jimthompson5910 (jim_thompson5910):

I'll get you started \[\large x + (x+1) \ge 46\] \[\large x + x+1 \ge 46\] \[\large 2x+1 \ge 46\] and I'll let you finish up since I have to get going. Good luck

OpenStudy (anonymous):

lol ok